OpenStudy (anonymous):

solve. 3 1/2 divided 7/8 +2 1/12 times 2/5 = a.2 13/30 b.1 5/9 c.1 1/12 d. 4 5/6

OpenStudy (anonymous):

= 7/2 * 8/7 + 25/12 * 2/5 = 56/14 + 50/60 = 28 /7 + 5/6 = 168/42 + 35/42 = 203/42 = 29 /6 = 4 5/6 d.